OpenStudy (anonymous):

write these fractions in order from least to greatest ?? 1/3 5/6 2/9 show work can sombodyhelp me

OpenStudy (anonymous):

put them all over a common denominator which is 18 1/3=6/18 5/6=15/18 2/9=4/18 and thus you can see that 2/9<1/3<5/6

OpenStudy (anonymous):

thankxx